# MeRF: Motivation-enhanced Reinforcement Finetuning for Large Reasoning Models

### v1

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful learn-to-reason paradigm for Large Language Models (LLMs) to tackle complex reasoning tasks. However, existing RLVR methods overlook one of the most distinctive capabilities of LLMs, their in-context learning ability, as prominently demonstrated by the success of Chain-of-Thought (CoT) prompting. This motivates us to explore how reinforcement learning can be effectively combined with in-context learning to better improve the reasoning capabilities of LLMs. In this paper, we introduce Motivation-enhanced Reinforcement Finetuning (MeRF), an intuitive yet effective method enhancing reinforcement learning of LLMs by involving “telling LLMs the rules of the game”. Specifically, MeRF directly injects the reward specification into the prompt, which serves as an in-context motivation for model to improve its responses with awareness of the optimization objective. This simple modification leverages the in-context learning ability of LLMs aligning generation with optimization, thereby incentivizing the model to generate desired outputs from both inner motivation and external reward. Empirical evaluations on the Knights and Knaves (K&K) logic puzzle benchmark demonstrate that MeRF achieves substantial performance gains over baselines. Moreover, ablation studies show that performance improves with greater consistency between the in-context motivation and the external reward function, while the model also demonstrates an ability to adapt to misleading motivations through reinforcement learning.
Figure 1: Validation Accuracy and Response Length of the RLVR and MeRF models on K&K puzzles benchmark. MeRF shows a significant improvement over RLVR in performance with in-context motivation in the reinforcement finetuning process. Meanwhile, the response length of MeRF models is comparable with the RLVR models, indicating our method of more efficient reasoning capability.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language understanding and generation tasks, such as instruction following (Ouyang et al., 2022; Zhou et al., 2023; Taori et al., 2023), code generation (Chen et al., 2021; Nijkamp et al., 2022; Zhuo et al., 2024), and medical diagnosis (Singhal et al., 2025; Zhang et al., 2023; Wang et al., 2023). To further improve the reasoning capabilities of LLMs, Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024; Team et al., 2025) has emerged as a promising alternative to conventional supervised fine-tuning approaches (Radford et al., 2018; Brown et al., 2020), as demonstrated by DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024). RLVR treats reasoning as a sequential decision-making process and optimizes models using objective reward signals that can be automatically verified, such as matching mathematical answers to ground truth (Lambert et al., 2024; He et al., 2024), or checking code correctness through unit tests (Le et al., 2022; Liu et al., 2023). This makes RLVR more scalable and robust compared to approaches like Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Rafailov et al., 2023), which rely on human preference annotations that can be noisy, inconsistent, and costly to obtain. By optimizing toward clearly defined success criteria, RLVR enables LLMs to iteratively refine their reasoning capabilities, achieving significant performance improvements on complex tasks.
@@
Extending the success of the conventional RL paradigm in gaming, Go, and robotics to LLMs, RLVR relys on external reward signals to guide the training process of LLMs (Zeng et al., 2025; Yu et al., 2025; Hu et al., 2025; Chen et al., 2025). However, RLVR in its current form does not fully exploit the in-context learning capabilities (Wei et al., 2022; Kojima et al., 2022) of LLMs, which is one of their most powerful features. In RLVR, the model receives feedback on the correctness of its outputs, but lacks explicit awareness of the optimization objectives during training. As the verifiable reward function has demonstrated the desired behavior of the model, which can described in natural language, and the in-context learning ability enables LLMs to learn from the given context, an intuitive question is why not tell the LLMs what is the expected behavior, or how is the their output get evaluated, in the reinforcement finetuning process? This is similar to how humans learn: when we are given a task, we often benefit from understanding the rules and objectives before we start working on it. This understanding helps us to align our efforts with the desired outcomes, making the learning process more efficient and effective.
@@
In this paper, we introduce Motivation-enhanced Reinforcement Finetuning (MeRF), a simple yet powerful method that injects the reward specification directly into the prompt, serving as an in-context motivation for model to improve its responses with awareness of the optimization objective. Unlike current RLVR approaches leaving the model blind to the optimization objective during generation, relying on the transcendent reward function to guide the training process, MeRF explicitly informs the model about the reward structure and what constitutes a good response with in-context motivation, incentivizing the model to generate desired outputs from both inner motivation and external reward, leading to more efficient reinforcement finetuning.
@@
  • We propose MeRF, a novel motivation-enhanced reinforcement finetuning method for LLMs, leveraging the in-context learning capabilities of LLMs in the reinforcement finetuning process for efficient improvement of reasoning capabilities.
@@
  • Extensive experiments on the K&K logic puzzle reasoning benchmark show that MeRF significantly outperforms RLVR baseline, validating its effectiveness and efficiency in improving reasoning capabilities of LLMs in complex tasks.
@@
  • We provide comprehensive analysis on the effectiveness of MeRF, including the impact of the consistency between the in-context motivation and the actual reward function, offering insights into the aligned in-context learning with reinforcement finetuning for self-evolving LLMs.
2 Preliminary
2.1 Reinforcement Learning with Verifiable Rewards (RLVR)
Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024) is a reinforcement learning paradigm for traning langauge models on tasks with verifiable outcomes, such as math problems or logic puzzles. The key idea is to use a reward function that can be automatically verified, allowing the model to learn from the ultimate correctness of its outputs. Recent works (Zeng et al., 2025; Yu et al., 2025) have shown that RLVR can significantly improve the reasoning capabilities of LLMs, with reasoning patterns emerged from the optimization for verifiable rewards.
2.2 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a reinforcement learning algorithm designed to optimize policies by leveraging group-wise relative preference information. As a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO foregoes the need for critic models and instead focuses on learning from relative comparisons of actions within groups, significantly enhancing training efficiency for reinforcement finetuning of LLMs. For each question $x$, GRPO samples a group of $G$ outputs $\{y_{i}\}_{i=1}^{G}$ from the policy $\pi_{\theta_{\text{old}}}(\cdot|x)$, and computes the advantage $A_{i}$ for each output $y_{i}$ based on the outcome reward $r_{i}$. The GRPO objective is defined as follows:
$\mathcal{L}_{\text{GRPO}}(\theta)$ $=\mathbb{E}\left[x\sim\mathcal{D};\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)\right]$ (1)
$=\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}A_{i},\text{clip}\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},1-\varepsilon,1+\varepsilon\right)A_{i}\right)-\beta\mathbb{D}_{KL}\left(\pi_{\theta}||\pi_{\text{ref}}\right)\right),$ (2)
where the advantage
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})}$
is computed by normalizing the outcome rewards of the group of ouputs and sets the advantage for all the tokens in the output $\{y_{i}\}_{i=1}^{G}$. This formulation enables GRPO to learn from relative preferences within each group without the need for a critic model, making it efficient and suitable for the implementation of RLVR to LLMs.
2.3 Knights and Knaves Logic Puzzles
The Knights and Knaves (K&K) (Xie et al., 2024) logic puzzle dataset is a widely used (Xie et al., 2025) benchmark for reinforcement finetuning for LLMs reasoning, which provides a well-structured difficulty level of the logic tasks and allows accurate and easy reward verification for RLVR. The controllable difficulty levels are achieved by varying the number of people in the logic task, the more people in the logic puzzle requiring LLMs more complex reasoning, and more steps to solve the task. The K&K dataset contains 7 different difficulty levels of logic puzzles, with 2 people as the easiest level and 8 people as the most difficult level. Here is an example of the K&K dataset with 3 people:
Figure 2: (1) Left: We validate models with motivation and without motivation for inference. Different from the startpoint model, RLVR model achieves a better performance in validation w/ motivation while the motivation is not involved in the training process. We find a connection between the in-context motivation validation and the model optimized through the reinforcement learning process, which is guided by the reward function described in the motivation. (2) Right: The motivation does not directly enhance in-context inference, but instead yields improvements progressively during the RLVR training.
3 Motivation-enhanced Reinforcement Finetuning
@@
In this section, we introduce the Motivation-enhanced Reinforcement Finetuning (MeRF) for efficient LLM reasoning, which enable the LLMs to be aware of the motivation of the reinforcement optimization and motivated by the specific objective in the verifiable reward space.
3.1 Motivation-enhanced Learning
Considering one’s learning process when facing a new task, one may be asked to do something without any information about the task, like "Please solve this math problem." In this case, one may try to solve the problem with the knowledge and experience one has, but it is very likely that one will not be able to solve it correctly, even though actually equipped with enough knowledge for solving it. The reason is that one does not know what is expected and how to do it correctly. One may be able to solve the problem after many trials and errors, but it is very inefficient and time-consuming. Intuitively, it would be much more efficient if one could be told the motivation of the task, like the goal of the task, how to do it correctly, and how would the output likely to be when it is correct. For example, when one is asked to solve a math problem, comparing to be told: "You should think step by step and give the correct answer in the <answer></answer> tag", it will be much more efficient if told: "You should think step by step and give the correct answer in the <answer></answer> tag, if you follow the steps as the demonstration, you will get a score of 1. You will be rewarded with an additional 1 point if you can give the correct answer in the <answer></answer> tag. Otherwise, you will get a score of 0 and no additional points. " This is a clear motivation for humans to solve math problems correctly, providing what is expected and how the metric is for the task.
Recently, RLVR has demonstrated using reinforcement learning to improve the performance of LLMs on various tasks (Team et al., 2025; Guo et al., 2025), expanding the success of reinforcement learning to LLMs while resembling the human learning process. However, the existing reinforcement learning methods for LLMs does not effectively utilize the in-context learning ability of LLMs, which is the key to the success of LLMs. For successful RL training, one of the most important components is a well-defined reward function, which shapes the expected behavior of the policy and guides the training process. In RLVR, the reward function quantifies the correctness of the model’s outputs, and sometimes with other metrics designed by humans for expected behavior. It is true that the reward function designer knows what is the expected behavior of the model, however, will it be possible for LLMs to perform better if they are aware of the motivation of the task, as humans do?
@@ 3.1 Motivation-enhanced Learning
First, we conduct an experiment, comparing the performance of the startpoint model and RLVR model with and without motivation in the inference time. The results are shown in Figure 2, where we find that the startpoint model does not benefit from the motivation in the inference time, while the RLVR model, trained with the reward function described in the motivation, achieves an improvement when validated with in-context motivation. Moreover, we also observe that the motivation does not directly enhance in-context inference, but instead yields improvements progressively during the RLVR training, indicating a connection between the in-context motivation and reinforcement learning process optimized by the transcendent reward function described in motivation. Motivated by this observation, we propose to inject motivation into the training process, leveraging the in-context ability of LLMs for more efficient reinforcement finetuning.
3.2 Motivation with Verified Reward
@@
Following the previous work (Xie et al., 2025) of RLVR, we utilize a verified reward function for the K&K logic puzzle dataset. The reward function contains two components: the Correctness Score and the Format Score. It is implemented by rule-based verification and capable of being described in natural language, enabling the motivation to be injected into the training process. Here is the System Prompt and Motivation for K&K puzzle:
@@
System Prompt and Motivation for K&K Puzzle <|im_start|>system You are a helpful assistant. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and<answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Now the user asks you to solve a logical reasoning problem. After thinking, when you finally reach a conclusion, clearly state the identity of each character within <answer> </answer> tags. i.e., <answer> (1) Zoey is a knight (2) … </answer>. You will get evaluated following Evaluation Scoring Rules: - Correctness Score: - If your final answer is correct, score 2 - If your answer is understandable but wrong, score -1.5 - If your answer is not parsable or incomplete, score -2 - Format Score: - If you follow the tag format exactly as above, score 1 - Otherwise, score -1 You will get the final score as their sum. Example: (1) The format follows the required structure: +1 (2) The final answer is correct: +2 (3) Total evaluation score: 3 Think carefully, follow the structure, and consider the evaluation rules.<|im_end|>
@@
By injecting the motivation into the training process, we enable the model to be aware of the motivation of the task, which describe the reward function of the RLVR pipeline. The motivation provides a clear specification of what is expected and how to do it correctly, aligning the generation with the transcendent optimization objective. This approach leverages the in-context learning ability of LLMs to improve their reasoning capabilities in a more efficient and effective manner.
@@ 4 Experiment
In this section, we present the experimental results of our proposed method MeRF for LLM reasoning. We compare the performance of MeRF with the RLVR baseline to demonstrate the effectiveness of motivation-enhanced learning in improving the reasoning capabilities of LLMs in the reinforcement finetuning process.
@@ 4.1 Experimental Setup
Models and RL algorithm
We conducted the motivation-enhanced experiments with Qwen2.5 series (Yang et al., 2024) as the startpoint. We used the Qwen2.5-7B-Instruct model as the startpoint model and use GRPO as the RL algorithm for reinforcement finetuning. Qwen2.5-7B-Instruct is a widely used LLM in the research community with a moderate size of 7 billion parameters. Qwen2.5-7B-Instruct is capable of powerful instruct-following for reasoning tasks, which is considered important for reinforcement finetuning of LLMs (Guo et al., 2025). GRPO is an effective and efficient reinforcement learning algorithm for LLMs finetuning, which is suitable for our experiments to demonstrate the effectiveness of MeRF and RLVR baseline in the reinforcement finetuning process without demanding excessive computational resources.
Dataset
We evaluate our method on the K&K logic puzzle dataset. The K&K dataset contains 7 different difficulty levels of logic puzzles, with ranging from 2 people to 8 people in the task. We utilize the 3 to 7 people puzzles for training, the corresponding test set for in-domain evaluation, and the 2 and 8 people puzzles for out-of-distribution (OOD) evaluation. There are 900 samples for training in each difficulty level and 100 samples for evaluation. The total number of samples in the training set is 4500, and 700 samples in the evaluation set.
Other Implementation Details
We implement our method with VeRL (Sheng et al., 2024) and Logic-RL (Xie et al., 2025) codebase. We use the Qwen2.5-7B-Instruct model as the startpoint and conduct the reinforcement finetuning with batch zise of 16, learning rate of 1e-6, 2 epochs of training. We use the rollout framework of Vllm (Kwon et al., 2023) for sampling the outputs of the model. The group size of GRPO is set to 8, clip range is set to 0.2, and the KL penalty is set to 0.001.
Table 1: Performance comparison across models on tasks with varying difficulty by number of people. MeRF demonstrates a significant improvement over the RLVR baseline in both in-domain and OOD scenarios. It is notable all the results are validated without in-context motivation.

### v2

# A Simple "Motivation" Can Enhance Reinforcement Finetuning of Large Reasoning Models

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful learn-to-reason paradigm for large reasoning models to tackle complex tasks. However, the current RLVR paradigm is still not efficient enough, as it works in a trial-and-error manner. To perform better, the model needs to explore the reward space by numerously generating responses and learn from fragmented reward signals, blind to the overall reward patterns. Fortunately, verifiable rewards make the natural language description of the reward function possible, and meanwhile, LLMs have demonstrated strong in-context learning ability. This motivates us to explore if large reasoning models can benefit from a motivation of the task, i.e., awareness of the reward function, during the reinforcement finetuning process, as we humans sometimes do when learning. In this paper, we introduce Motivation-enhanced Reinforcement Finetuning (MeRF), an intuitive yet effective method enhancing reinforcement finetuning of LLMs by involving “telling LLMs rules of the game”. Specifically, MeRF directly injects the reward specification into the prompt, which serves as an in-context motivation for the model to be aware of the optimization objective. This simple modification leverages the in-context learning ability of LLMs, aligning generation with optimization, thereby incentivizing the model to generate desired outputs from both inner motivation and external reward. Empirical evaluations demonstrate that MeRF achieves substantial performance gains over the RLVR baseline. Moreover, ablation studies show that MeRF performs better with greater consistency between the in-context motivation and the external reward function, while the model also demonstrates an ability to adapt to misleading motivations through reinforcement finetuning.
Figure 1: Validation Accuracy of MeRF and RLVR baseline on K&K Logic Puzzles. Simply “telling LLMs rules of the game” with in-context motivation during training, MeRF significantly outperforms the RLVR baseline with better efficiency, demonstrating the effectiveness of leveraging in-context motivation for more efficient RL training.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language understanding and generation tasks, such as instruction following (Ouyang et al., 2022; Zhou et al., 2023; Taori et al., 2023), code generation (Chen et al., 2021; Nijkamp et al., 2022; Zhuo et al., 2024), and medical diagnosis (Singhal et al., 2025; Zhang et al., 2023; Wang et al., 2023). To further improve the reasoning capabilities of LLMs, Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024; Team et al., 2025) has emerged as a promising alternative to conventional supervised fine-tuning approaches (Radford et al., 2018; Brown et al., 2020), as demonstrated by DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024). RLVR treats reasoning as a sequential decision-making process and optimizes models using objective reward signals that can be automatically verified with explicit rules, such as matching mathematical answers to ground truth (Lambert et al., 2024; He et al., 2024), or checking code correctness through unit tests (Le et al., 2022; Liu et al., 2023). By optimizing models toward meeting the clearly defined success criteria, RLVR enables LLMs to iteratively refine their reasoning capabilities, achieving significant performance improvements on complex tasks.
Following and extending the success of the conventional RL paradigm in gaming, Go, and robotics to the field of language models, RLVR relies on external reward signals to guide the training process of LLMs (Zeng et al., 2025; Yu et al., 2025; Hu et al., 2025; Chen et al., 2025). However, in the same way as conventional reinforcement learning, RLVR is not very efficient, as it works in a trial-and-error manner, where the model can suffer from the sparse reward space and needs to learn the patterns of the task and reward function by repeatedly collecting and comparing its outputs and corresponding rewards. It requires a larger amount of training data and computational resources for LLMs to learn from and generate enough responses to reach the expected behavior and then get positive feedback. As we continue expecting LLMs to solve more and more complex tasks, the reward function can be more sophisticated and the expected behavior can be harder to reach, making it even more important to improve the efficiency of the current RLVR paradigm.
@@
In RLVR, the model receives feedback on the correctness of its outputs, but lacks explicit awareness of the optimization objectives during training. As the verifiable reward function (with explicit rules) has demonstrated the desired behavior of the model, which can be described in natural language, and the in-context learning ability enables LLMs to learn from the given context, an intuitive question is why not tell the LLMs what is the expected behavior, or how is their output get evaluated, during the training? This is similar to how humans learn: when we have a task, we often benefit from understanding the rules and objectives before we start working on it. This understanding helps us to align our efforts with the desired outcomes, resulting in more efficient and effective learning.
@@
In this paper, we introduce Motivation-enhanced Reinforcement Finetuning (MeRF), a simple yet powerful method that injects the reward specification directly into the prompt, serving as an in-context motivation for the model to be aware of the optimization objective. Unlike current RLVR paradigm leaving the model blind to the optimization objective during generation, relying on the transcendent reward function to guide the training process, MeRF explicitly informs the model about the reward structure and what constitutes a good response with in-context motivation, incentivizing the model to generate desired outputs from both inner motivation and external reward, leading to more efficient reinforcement finetuning as shown in Figure 1.
@@
  • We propose MeRF, a novel motivation-enhanced reinforcement finetuning method for LLMs, enabling the model to be aware of the optimization objective by in-context motivation, to achieve more efficient and effective reinforcement finetuning of large reasoning models.
@@
  • Extensive experiments on the reasoning benchmarks: K&K Logic Puzzles, AIME24&25, AMC23, MATH500, and CountDown, show that MeRF significantly outperforms the RLVR baseline, validating its effectiveness and efficiency in improving reasoning capabilities in complex tasks.
@@
  • We provide a comprehensive analysis on the effectiveness of MeRF, including the impact of the consistency between the in-context motivation and the actual reward function, offering insights into the aligned in-context learning with reinforcement finetuning for self-evolving LLMs.
2 Method
@@
In this section, we introduce the Motivation-enhanced Reinforcement Finetuning (MeRF) for more efficient reinforcement learning with verifiable rewards of large reasoning models, which enables LLMs to be aware of the objective of the task in the reward space by in-context motivation.
Figure 2: (Left) Illustration of the RLVR pipeline and the in-context motivation introduced by MeRF. Compared to the indirect way (reward samples generated and through parameter updates) to learn the reward patterns, MeRF enables the model to be aware of the overall reward space by in-context motivation. (Right) We validate models in two settings: w/ motivation and w/o motivation in the prompt. Different from the base model, the RLVR model achieves a slightly better performance in validation w/ motivation than w/o motivation after the RLVR training, even while the motivation is not involved in the training process, indicating a connection between the in-context motivation validation and the RLVR training guided by the reward function (as the motivation describes).
2.1 Motivation-enhanced Reinforcement Finetuning
Reinforcement learning of LLMs performs in a similar way as the human learning process, gaining improvements by learning from the feedback from the environment. The feedback is critical for the learning process, for it determines the direction of the optimization, which is often in the form of rewards, reflecting how well the model performs on the task. However, it is not easy for the model to learn the reward patterns before exploring the reward space extensively, especially when the reward space is sparse and the expected behavior is hard to reach. It can be inefficient and computationally expensive for LLMs, and at the beginning, most of the responses generated are bad and non-rewarded. Models work in a blind way, i.e., randomly respond to the question, and somehow the precious fragments of sparse positive rewards help the model to learn the reward patterns and expected behavior. This RLVR paradigm is not efficient enough and sometimes even fails to work, when the current model can hardly generate any good responses to get better rewards.
The cause of the problem is: the model is optimized in an indirect way with a black-box manner: the model is unaware of the overall optimization objective of training during the generation, gaining reward signal information by fragments (reward samples by the exploration of the model), and through the parameter updates (policy gradient). The reward information can be local and one-sided when exploration fails to sufficiently cover the overall reward space (which is challenging), preventing the model from reaching the global optimum and leading instead to reward hacking toward local optima. As shown in Figure 2 (Left), in current RLVR pipeline, the model can’t learn how to achieve the +10 reward if none of the generated responses can reach it, trapped in a paradoxical learning situation: you need to learn something that you don’t know how to do, or even don’t know it exists.
Fortunately, the problem can be alleviated in RLVR and LLMs. In the RLVR pipeline, the reward is verifiable with explicit rules, which means the reward function can be described by natural language, and meanwhile, LLMs have demonstrated strong in-context learning ability to learn from the given context. We propose to improve the information flow of reward signals in a more direct way, by in-context motivation, i.e., language description of the reward function, in the training process, to make the model aware of the optimization objective.
@@ 2.1 Motivation-enhanced Reinforcement Finetuning
In Figure 2 (Right), we conduct an experiment, comparing the performance of the Base model and RLVR model with and without motivation in the inference time, where we find that the Base Model does not benefit from the motivation in the inference time, while the RLVR model, trained with the reward function described in the motivation, achieves an improvement when validated with in-context motivation. Moreover, when we involve the motivation during the training process (MeRF), we find a significant performance improvement over the RLVR baseline, demonstrating the effectiveness of MeRF in improving the efficiency and effectiveness of reinforcement finetuning of LLMs.
2.2 Motivation with Verifiable Reward
@@
Following the previous work (Xie et al., 2025) of RLVR, we utilize a verifiable reward function for the K&K Logic Puzzle dataset and demonstrate how the motivation is designed based on the reward function. The reward function contains two components: (i) Correctness Score and (ii) Format Score. It is implemented by rule-based verification and capable of being described in natural language, enabling the motivation to be injected into the training process. Here is the System Prompt and Motivation for K&K puzzle:
@@
System Prompt and Motivation for K&K Puzzle <|im_start|>system You are a helpful assistant. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and<answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Now the user asks you to solve a logical reasoning problem. After thinking, when you finally reach a conclusion, clearly state the identity of each character within <answer> </answer> tags. i.e., <answer> (1) Zoey is a knight (2) … </answer>. You will get evaluated following Evaluation Scoring Rules:
- Correctness Score:
 - If your final answer is correct, score 2
 - If your answer is understandable but wrong, score -1.5
 - If your answer is not parsable or incomplete, score -2
- Format Score:
 - If you follow the tag format exactly as above, score 1
 - Otherwise, score -1
You will get the final score as their sum. Example:
(1) The format follows the required structure: +1
(2) The final answer is correct: +2
(3) Total evaluation score: 3
Think carefully, follow the structure, and consider the evaluation rules.
<|im_end|>
@@
By injecting the motivation into the training process, we enable the model to be aware of the motivation of the task, which describes the reward function of the RLVR pipeline. The motivation provides a clear specification of what is expected and how to do it correctly, aligning the generation with the transcendent optimization objective. This approach leverages the in-context learning ability of LLMs to improve their reasoning capabilities in a more efficient and effective manner.
@@ 3 Experiment
In this section, we present the experimental results of our proposed method MeRF. We compare the performance of MeRF with the RLVR baseline to demonstrate the effectiveness of MeRF in improving the reasoning capabilities of LLMs in the reinforcement finetuning process.
@@ 3.1 Experimental Setup
Models and RL Algorithm.
We conducted the experiments with Qwen2.5 series (Yang et al., 2024) and DeepSeek-R1-Distill series (Guo et al., 2025), including: Qwen2.5-7B-Base, Qwen2.5-7B-Instruct, Qwen2.5-7B-Instruct-1M, Qwen2.5-14B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Llama-8B. The models are across different model sizes, model families, and instruction-tuning stages, allowing us to comprehensively evaluate the effectiveness of MeRF in enhancing reinforcement finetuning of LLMs. We use GRPO as the RL algorithm for reinforcement finetuning. GRPO is an effective and efficient reinforcement learning algorithm for LLMs finetuning, which is suitable for our experiments to demonstrate the effectiveness of MeRF and RLVR baseline in the reinforcement finetuning process without demanding excessive computational resources.
Dataset.
We evaluate MeRF on K&K (Logic Puzzles), MATH benchmarks: AIME24&25, AMC23, MATH500 (Lightman et al., 2023), and CountDown (Number Game) (Pan et al., 2025). The K&K dataset contains 7 different difficulty levels of logic puzzles, ranging from 2 people to 8 people in the task. We utilize the 3 to 7 people puzzles for training, the corresponding test set for in-domain evaluation, and the 2 and 8 people puzzles for out-of-distribution (OOD) evaluation. There are 900 samples for training in each difficulty level and 100 samples for evaluation. The total number of samples in the training set is 4500, and 700 samples for evaluation. For MATH benchmarks, we follow the previous work (Yu et al., 2025), using a subset of the DAPO-Math-17K dataset for training and evaluating on AIME24&25, AMC23, and MATH500. We modify the original prompt of the training data for motivation design. For CountDown, we follow the previous work (Pan et al., 2025), using the same training and evaluation set. CountDown is a number game, where the model needs to use the given numbers and arithmetic operations to reach the target number. The given numbers can only be used once, and the number of given numbers is 3 or 4 in our experiments.
Table 1: Performance comparison across models on tasks with varying difficulty by number of people of K&K Puzzles. MeRF demonstrates a significant improvement over the RLVR baseline in both in-domain and OOD scenarios. Notably, all the results are validated without in-context motivation.